# OntoURL: A Benchmark for Evaluating Large Language Models on Symbolic Ontological Understanding, Reasoning and Learning

### v2

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capabilities across a range of natural language processing tasks, yet their ability to process structured symbolic knowledge remains underexplored. To address this gap, we propose a taxonomy of LLMs’ ontological capabilities and introduce OntoURL, the first comprehensive benchmark designed to systematically evaluate LLMs’ proficiency in handling ontologies—formal, symbolic representations of domain knowledge through concepts, relationships, and instances. Based on the proposed taxonomy, OntoURL systematically assesses three dimensions: understanding, reasoning, and learning through 15 distinct tasks comprising 58,981 questions derived from 40 ontologies across 8 domains. Experiments with 20 open-source LLMs reveal significant performance differences across models, tasks, and domains, with current LLMs showing proficiency in understanding ontological knowledge but substantial weaknesses in reasoning and learning tasks. These findings highlight fundamental limitations in LLMs’ capability to process symbolic knowledge and establish OntoURL as a critical benchmark for advancing the integration of LLMs with formal knowledge representations.
@@ 1 Introduction
Ontologies play a foundational role in encoding structured domain knowledge among the most prominent symbolic frameworks, particularly in fields such as finance, medicine, science and legal domain. They provide a formal structure through well-defined concepts (classes), relationships (e.g., hierarchies and semantic connections), and instances (individuals) gruber1993translation ; noy2001ontology ; mcguinness2004owl . In recent years, as Large language models (LLMs) have significantly revolutionized numerous fields with their remarkable capabilities in various tasks openai2023gpt4 ; wei-et-al-2022-chain , the interaction between ontologies and the knowledge-rich LLMs has sparked significant interest, raising research into ontology-related tasks such as leveraging LLMs for ontology creation.
@@
However, whether LLMs can truly comprehend and manipulate structured symbolic knowledge remains a subject of intense debate tang2023large ; pavlick2023symbols ; yan2024large ; saba2024reinterpreting . This discussion centers on whether statistical pattern recognition can replicate the symbolic representations and logical structures traditionally managed by classical knowledge representation systems. Therefore, a critical yet underexplored question arises for ontology practitioners and researchers: to what extent can LLMs understand, utilize, and construct ontologies? While several datasets have been developed for ontology-related tasks (badie2017formal, ; peng-etal-2022-copen, ; wu-etal-2023-plms, ; bombieri2024llms, ; qin2024comprehensive, ; song2025peeled, ; he2023language, ; sun2024large, ; kang2024researcharena, ; jiang2025hibench, ; e2eol2024lo, ; li2024ontology, ), these works typically focus on one or two isolated ontology aspects and are rarely designed specifically for evaluating LLMs. Furthermore, there is an absence of a comprehensive taxonomy that systematically categorizes the ontological capabilities required across domains and tasks, thus hindering the reliable evaluation of LLMs capabilities on ontology.
@@
To fill in this gap, we investigate three key research questions spanning three dimensions: (1) can LLMs accurately memorize the fine-grained details inherent in ontologies, including concepts, hierarchical relationships, properties, and instances? (2) Can LLMs perform robust reasoning over ontologies, such as transitive superclass inference or description logic reasoning? (3) Can LLMs autonomously construct ontologies based on their rich knowledge, such as tasks like ontology hierarchy construction?
@@
We first propose a taxonomy of ontological capabilities for LLMs and then introduce OntoURL, the first comprehensive benchmark for evaluating LLMs’ abilities on ontologies. OntoURL consists of 58,981 questions derived from 40 ontologies, encompassing 15 tasks across 8 domains. These tasks are meticulously crafted to assess LLMs’ proficiency in three crucial dimensions: understanding, reasoning, and learning. By evaluating 20 open-source LLMs, we gain critical insights into their strengths and limitations on handling structured symbolic knowledge. Our primary contributions are summarized as follows:
@@
  • We present a taxonomy of LLM ontological capabilities in three dimensions: understanding, reasoning, and learning, providing a systematic framework for analyzing LLMs’ interactions with structured symbolic knowledge.
@@
  • We introduce OntoURL, a benchmark comprising 58,981 questions derived from 40 ontologies, covering 15 tasks across 8 domains. Our benchmark enables rigorous evaluation on LLMs’ abilities across multiple dimensions, such as conceptual understanding, structural knowledge, logical reasoning, structure construction, and concept alignment.
@@
  • We conduct in-depth pilot studies on 20 LLMs and provide comprehensive analyses across model scales, task levels, and specific domains. All code for benchmark construction and experiments, as well as the benchmark dataset, are publicly available (Appendix A).
2 Background
@@
Ontologies provide formal, logic-based representations of domain knowledge, organizing concepts, properties, instances, and their interrelations in a structured, symbolic form gruber1993translation (Appendix C.1). Unlike taxonomies or controlled vocabularies which focus primarily on hierarchical groupings, or knowledge graphs which typically emphasize instance-level assertions, ontologies encode domain semantics through explicit axioms and logical constraints (further comparison in Appendix B.3).
As a formal knowledge representation system, ontologies are typically formalized in Description Logic krötzsch2013descriptionlogicprimer —a family of formal languages designed for representing and reasoning about knowledge. Ontologies comprise three fundamental components: Terminological Box (TBox), containing class hierarchies and definitions; Assertional Box (ABox), capturing assertions about individuals; and Role Box (RBox), defining relationships and properties among entities. Figure 1 illustrates these components in a conference ontology excerpt, where “Chair” and its superclass “Committee Members” represent TBox elements, “Mary” as an instance of “Chair” forms part of the ABox, and “has_authors” is a relation between “Review” and “Review Expertise” belongs to the RBox.
By offering a standardized vocabulary with formal semantics, ontologies support semantic interoperability, knowledge integration, and automated reasoning staab2013handbook . Most ontological resources are created by domain experts, such as Gene Ontology ashburner2000gene , Plant Ontology plant2002plant , and LKIF Core Legal Ontology hoekstra2007lkif . Appendix C.4 summarizes the 40 ontologies used in our benchmark, and Appendix C.3 introduces the language for describing these ontologies: RDF (a foundational data model) and OWL (the core ontology language). As ontologies have been central to symbolic AI approaches for decades, understanding and leveraging such structured symbolic knowledge are essential for LLMs.
@@
Previous work has primarily focused on conceptual understanding, using probing techniques to examine how LLMs memorize and retrieve class-level knowledge badie2017formal ; peng-etal-2022-copen ; sahu2022unpacking ; patel2022mapping ; wu-etal-2023-plms ; shani-etal-2023-towards ; jang-lukasiewicz-2023-improving ; mitchell2023debate ; jin2024exploring ; song2025peeled , and structural knowledge he2023language ; mruthyunjaya2023rethinking ; park2024geometry ; jackermeier2024dual ; zhang-etal-2025-neural .
Beyond basic understanding tasks, some works perform specialized forms of deductive logic reasoning on ontologies. Rule-based ontology reasoners support a wide range of inference tasks, including validating ontology coherence, deriving complete class hierarchies, assigning individuals to their most specific types, inferring property relationships, and executing queries to retrieve relevant classes or individuals tsarkov2006fact ; mendez2009reintroducing ; kazakov2012elk ; sertkaya2013elephant ; glimm2014hermit ; ceylan2015bayesian ; bobillo2016fuzzy ; fernandes2018graph ; balhoff2018arachne . Recently, a few studies have begun to explore using language models for ontology reasoning, particularly within the framework of Description Logic he2023language ; wang2024can .
Another active research direction involves learning and constructing ontological structures. Traditional approaches employed statistical term extraction and pattern-based methods to identify candidate concepts and taxonomic relations hearst1998automated ; kietz2000method ; maedche2001ontology ; xu2002domain ; alfonseca2002unsupervised ; lonsdale2002peppering ; khan2002ontology ; biemann2005ontology ; asim2018survey ; xu2019automatic ; konys2019knowledge . More recent efforts leverage pretrained language models, enabling more sophisticated concept extraction and hierarchy learning babaei2023llms4ol ; neuhaus2023ontologies ; e2eol2024lo .
@@
To evaluate these diverse ontology-related tasks, researchers have developed various benchmarks focusing on specific capabilities. These include evaluations for conceptual knowledge badie2017formal ; peng-etal-2022-copen ; wu-etal-2023-plms ; bombieri2024llms ; qin2024comprehensive ; song2025peeled , hierarchical knowledge he2023language ; sun2024large ; kang2024researcharena ; jiang2025hibench , ontology reasoning he2023language ; wang2024can , ontology matching shvaiko2011ontology ; kolyvakis2018deepalignment ; kolyvakis2018biomedical ; iyer-etal-2021-veealign ; ibrahim2023toward , and ontology learning jiang2010crctol ; babaei2023llms4ol ; e2eol2024lo ; li2024ontology . However, most existing datasets and benchmarks focus on only one or two specific aspects of ontologies, and few are designed specifically for LLMs with appropriate question-answering or generation formats. This limitation underscores the need for a comprehensive benchmark that covers a wide range of ontologies, domains, and tasks.
@@ 3.1 Design Principle
OntoURL is designed as a comprehensive evaluation benchmark to systematically assess the multi-dimensional capabilities of LLMs within domain-specific ontologies. It serves two primary purposes: supporting ontology practitioners in selecting appropriate LLMs for ontology-related applications, and providing LLM researchers with insights into models’ conceptual, hierarchical, reasoning and generative capabilities in ontological contexts.
While considerable research has explored interactions between LLMs and ontologies, few studies have provided a systematic classification of the underlying capabilities. Drawing inspiration from Bloom’s Taxonomy of educational objectives bloom1956taxonomy , we introduce a three-level taxonomy for ontological capabilities for LLMs—understanding, reasoning and learning (Figure 2).
Figure 2: The taxonomy of LLM ontological capabilities, inspired by Bloom’s taxonomy. Each capability is positioned within a triangular structure and briefly explained on the left.
@@
This capability represents the most fundamental ontological level and is thus placed at the base of the triangle in Figure 2. It encompasses the memorization, recall, and comprehension of explicitly defined ontology knowledge. For example, retrieving the definition of the concept “calyx”in the Plant Ontology, identifying its superclass and subsumption relationships, and recognizing its associated properties and instances.
@@
[…]
@@
This capability represents the highest level in our taxonomy and is placed at the top of the triangle in Figure 2. It primarily concerns the process of constructing ontologies. Traditional ontology learning tasks have largely focused on generating hierarchical structures, while often neglecting other essential structural components like properties and instances. Therefore, we propose that ontological learning should encompass multiple dimensions: the generation of class definitions, the construction of class hierarchies, and the integration of properties and their constraints. In addition, ontology alignment—ensuring consistency across multiple ontologies by identifying and mapping equivalent concepts—is a critical aspect of ontology construction and is thus also considered part of this capability.
@@ 3.2 Data Collection and Processing
OntoURL draws on 40 expert-created, open-source ontologies spanning a broad range of 8 different domains, including (1) sciences; (2) health and medicine; (3) business and finance; (4) earth and environment; (5) arts, media and entertainment; (6) food and agriculture; (7) human and society; and (8) the legal domain. All ontologies are provided in RDF miller1998introduction or OWL mcguinness2004owl format. Detailed information of these ontologies can be found in Appendix C.4. While open-domain ontologies such as DBpedia auer2007dbpedia are available, we focus on domain-specific ontologies due to their greater depth, consistency, and formal structure (see Appendix B.4 for discussion). In cases where multiple ontologies exist within the same domain, we address their heterogeneity by designing prompts tailored to each ontology (more discussion in Appendix B.5).
@@
After collecting the ontologies, we apply a four-step pipeline to create multiple-choice, true/false question and open-ended questions, as illustrated in Figure 3. We provide more details and implementations of specific tasks in Appendix D.
First, we extract task-relevant entities, including classes, instances, properties and their associated semantic details (e.g., definitions, relationships, and range/domains) from each ontology. The sub-ontology of “Basenji” in the first step of Figure 3 presents the most involved elements in this extraction process. In particular, for reasoning tasks, an ontology reasoner is required to derive implicit relations (e.g. the relation between “Basenji” and “Canid”).
@@
After question generation, we apply several filtering and balancing strategies to ensure quality and diversity: (a) To avoid over-representing abstract concepts, we assign sampling probabilities based on class depth (distance from the root) and sample data according to these probabilities. (b) Questions for same classes are de-duplicated. For instance, as shown in Figure 3, a class like “Dog” has multiple subclasses (e.g., “Basenji”, “Leonberger”), but only one question about the subsumptions of "Dog" will be retained. (c) We limit the number of questions per class (they may from different ontologies) to a maximum of five per task. (d) We set the total number of questions per ontology-task pair at 500 to prevent any single ontology from disproportionately influencing the evaluation results.
Finally, we perform an human verification to ensure the data quality, focusing on three key dimensions: (1) Syntax, to verify fluentness and grammatical correctness; (2) Semantics, to confirm the correctness of the answer; and (3) Quality (for multiple-choice questions), to assess whether the distractors are well-designed and appropriately challenging. Each expert reviewed 20% of the data for each task, with a 5% overlap between annotators to allow cross-verification of annotation consistency. Expert review confirms the high quality of the benchmark, with over 95% of questions rated as syntactically and semantically correct, and inter-annotator agreement exceeding 0.85 (Fleiss’ $\kappa$ landis1977measurement ) across evaluation dimensions (detailed guidelines and results are provided in Appendix E).

### v3

@@ Abstract
Large language models have demonstrated remarkable capabilities across a wide range of tasks, yet their ability to process structured symbolic knowledge remains underexplored. To address this gap, we propose a taxonomy of ontological capabilities and introduce OntoURL, the first comprehensive benchmark designed to systematically evaluate LLMs’ capabilities in handling ontologies—formal and symbolic representations of domain knowledge. Based on the proposed taxonomy, OntoURL systematically assesses three dimensions: understanding, reasoning, and learning through 15 distinct tasks comprising 57,303 questions derived from 40 ontologies across 8 domains. Experiments with 20 open-source LLMs reveal significant performance differences across models, tasks, and domains, with current LLMs showing capabilities in understanding ontological knowledge but weaknesses in reasoning and learning tasks. Further experiments with few-shot and chain-of-thought prompting illustrate how different prompting strategies affect model performance. Additionally, a human evaluation reveals that LLMs outperform humans in understanding and reasoning tasks but fall short in most learning tasks. These findings highlight both the potential and limitations of LLMs in processing symbolic knowledge and establish OntoURL as a critical benchmark for advancing the integration of LLMs with formal knowledge representations.
@@ 1 Introduction
Ontologies play a foundational role in encoding structured domain knowledge among the most prominent symbolic frameworks, particularly in fields such as finance, the sciences, and law. They provide a formal structure through well-defined concepts (classes), relationships (e.g., hierarchies and semantic connections), and instances (individuals) Gruber (1993); Noy et al. (2001); McGuinness et al. (2004). In recent years, as Large language models (LLMs) have transformed numerous fields with their remarkable capabilities in various tasks Wei et al. (2022); OpenAI (2023), the interaction between ontologies and knowledge-rich LLMs has sparked significant interest, raising research into ontology-related tasks such as leveraging LLMs for ontology creation.
@@
However, whether LLMs can truly comprehend and manipulate structured symbolic knowledge remains a subject of intense debate Tang et al. (2023); Pavlick (2023); Yan et al. (2024); Saba (2024). This discussion centers on whether statistical pattern recognition can replicate the symbolic representations and logical structures traditionally managed by classical knowledge representation systems. Therefore, a critical yet underexplored question arises for ontology practitioners and researchers: to what extent can LLMs understand, utilize, and construct ontologies? While several datasets have been developed for ontology-related tasks (Wu et al., 2023; Bombieri et al., 2024; Qin et al., 2024; Song et al., 2025; He et al., 2023; Sun et al., 2024; Jiang et al., 2025; Lo et al., 2024; Li et al., 2024), these works typically focus on one or two isolated ontology aspects and are rarely designed specifically for evaluating LLMs. Furthermore, there is an absence of a comprehensive taxonomy that systematically categorizes the ontological capabilities required across domains and tasks, thus hindering the reliable evaluation of LLMs capabilities on ontology.
@@
To fill this gap, we investigate three key research questions: (1) can LLMs accurately memorize the fine-grained details inherent in ontologies, including concepts, hierarchical relationships, properties, and instances? (2) Can LLMs perform robust reasoning over ontologies, such as transitive superclass inference or description logic reasoning? (3) Can LLMs autonomously construct ontologies based on their rich knowledge, such as ontology hierarchy construction?
@@
We first propose a taxonomy of ontological capabilities for LLMs and then introduce OntoURL, the first comprehensive benchmark for evaluating LLMs’ abilities on ontologies. OntoURL consists of 57,303 questions derived from 40 ontologies, encompassing 15 tasks across 8 domains. These tasks meticulously assess LLMs’ proficiency in three crucial dimensions: understanding, reasoning, and learning. By evaluating 20 open-source LLMs, we gain critical insights into their strengths and limitations on handling structured symbolic knowledge. Our primary contributions are summarized as follows:
@@
  • We present a taxonomy of ontological capabilities in three dimensions: understanding, reasoning, and learning, providing a systematic framework for analyzing LLMs’ interactions with structured symbolic knowledge.
@@
  • We introduce OntoURL, a benchmark comprising 57,303 questions derived from 40 ontologies, covering 15 tasks across 8 domains. Our benchmark enables rigorous evaluation on LLMs’ abilities across multiple dimensions, such as conceptual understanding, logical reasoning, structure construction and alignment.
@@
  • We conduct in-depth pilot studies on 20 LLMs and provide comprehensive analyses across model scales, task levels, and specific domains. Some robust prompting strategies, such as few-shot and chain-of-thought prompting are also evaluated. An additional experiment on human performance are also provided. All code and data are available (Appendix A).
2 Background and Related Work
@@
Ontologies provide formal, logic-based representations of domain knowledge, organizing concepts, properties, instances, and their interrelations in a structured, symbolic form Gruber (1993). Unlike taxonomies or controlled vocabularies which focus primarily on hierarchical groupings, or knowledge graphs which typically emphasize instance-level assertions, ontologies encode domain semantics through explicit axioms and logical constraints.
As a formal knowledge representation system, ontologies are typically formalized in Description Logic Krötzsch et al. (2013)—a family of formal languages designed for representing and reasoning about knowledge. Ontologies comprise three fundamental components: Terminological Box (TBox), containing class hierarchies and definitions; Assertional Box (ABox), capturing assertions about individuals; and Role Box (RBox), defining relationships and properties among entities. Figure 1 illustrates these components in a conference ontology excerpt, where “Chair” and its superclass “Committee Members” represent TBox elements, “Mary” as an instance of “Chair” forms part of the ABox, and “has_authors” is a relation between “Review” and “Review Expertise” constitutes an RBox relation.
By offering a standardized vocabulary with formal semantics, ontologies support semantic interoperability, knowledge integration, and automated reasoning Staab and Studer (2013). Most ontological resources are created by domain experts, such as Gene Ontology Ashburner et al. (2000), Plant Ontology Consortium (2002), and LKIF Core Legal Ontology Hoekstra et al. (2007). As ontologies have been central to symbolic AI approaches for decades, understanding and leveraging such structured symbolic knowledge are essential for LLMs.
@@
Previous work has primarily focused on conceptual understanding, using probing techniques to examine how LLMs memorize and retrieve class-level knowledge Badie (2017); Peng et al. (2022); Sahu et al. (2022); Patel and Pavlick (2022); Wu et al. (2023); Shani et al. (2023); Jang and Lukasiewicz (2023); Mitchell and Krakauer (2023); Jin et al. (2024); Song et al. (2025), and structural knowledge He et al. (2023); Mruthyunjaya et al. (2023); Park et al. (2024); Jackermeier et al. (2024); Zhang et al. (2025).
Beyond basic understanding tasks, some works perform specialized forms of deductive logic reasoning on ontologies. Rule-based ontology reasoners support a wide range of inference tasks, including validating ontology coherence, deriving complete class hierarchies, assigning individuals to their most specific types, inferring property relationships, and executing queries to retrieve relevant classes or individuals Tsarkov and Horrocks (2006); Mendez and Suntisrivaraporn (2009); Kazakov et al. (2012); Sertkaya (2013); Glimm et al. (2014); Ceylan et al. (2015); Bobillo and Straccia (2016); Fernandes et al. (2018); Balhoff et al. (2018). Recently, a few studies have begun to explore using language models for ontology reasoning, particularly within the framework of Description Logic He et al. (2023); Wang et al. (2024).
Another active research direction involves learning and constructing ontological structures. Traditional approaches employed statistical term extraction and pattern-based methods to identify candidate concepts and taxonomic relations Hearst (1998); Kietz et al. (2000); Maedche and Staab (2001); Xu et al. (2002); Alfonseca and Manandhar (2002); Lonsdale et al. (2002); Khan and Luo (2002); Biemann (2005); Asim et al. (2018); Xu et al. (2019); Konys (2019). More recent efforts leverage pretrained language models, enabling more sophisticated concept extraction and hierarchy learning Babaei Giglou et al. (2023); Neuhaus (2023); Lo et al. (2024).
@@
To evaluate these diverse ontology-related tasks, researchers have developed various benchmarks focusing on specific capabilities. These include evaluations for conceptual knowledge Badie (2017); Peng et al. (2022); Wu et al. (2023); Bombieri et al. (2024); Qin et al. (2024); Song et al. (2025), hierarchical knowledge He et al. (2023); Sun et al. (2024); Kang and Xiong (2024); Jiang et al. (2025), ontology reasoning He et al. (2023); Wang et al. (2024), ontology matching Shvaiko and Euzenat (2011); Kolyvakis et al. (2018a, b); Iyer et al. (2021); Ibrahim et al. (2023), and ontology learning Jiang and Tan (2010); Babaei Giglou et al. (2023); Lo et al. (2024); Li et al. (2024). However, most existing datasets and benchmarks focus on only one or two specific aspects of ontologies, and few are designed specifically for LLMs with appropriate question-answering or generation formats. This limitation underscores the need for a comprehensive benchmark that covers a wide range of ontologies, domains, and tasks.
@@ 3.1 Design Principle
OntoURL is designed as an evaluation benchmark to systematically assess the multi-dimensional capabilities of LLMs within domain-specific ontologies. It serves two primary purposes: supporting ontology practitioners in selecting appropriate LLMs for ontology-related applications, and providing LLM researchers with insights into models’ conceptual, hierarchical, reasoning and generative capabilities in ontological contexts.
While considerable research has explored interactions between LLMs and ontologies, few studies have provided a systematic classification of the underlying capabilities. Drawing inspiration from Bloom’s Taxonomy of educational objectives Bloom et al. (1956), we introduce a three-level taxonomy for ontological capabilities for LLMs—understanding, reasoning and learning (Figure 2).
Figure 2: The taxonomy of LLM ontological capabilities, inspired by Bloom’s taxonomy. Each capability is positioned within a triangular structure and briefly explained on the right.
@@
This represents the most fundamental ontological level and is thus placed at the base of the triangle in Figure 2. It encompasses the memorization, recall, and comprehension of explicitly defined ontology knowledge. For example, retrieving the definition of the concept “calyx” in the Plant Ontology, identifying its superclass and subsumption relationships, and recognizing its associated properties and instances.
@@
This capability builds upon ontological understanding and is positioned above it in the triangle of Figure 2. It involves inferring implicit knowledge that is not explicitly defined within an ontology. Structured ontologies often encode rich hierarchical relationships from which additional facts can be logically deduced. For example, the Plant Ontology states that “testa” is a subclass of “seed coat” ($\texttt{testa}\sqsubseteq\texttt{seed\ coat}$), “seed coat” is a subclass of “seed” (seed $\texttt{seed\ coat}\sqsubseteq\texttt${seed}), and “seed” is a subclass of “plant embryo” ($\texttt{seed}\sqsubseteq\texttt{plant\ embryo}$). From these axioms, it follows that “testa” is also a subclass of “plant embryo” ($\texttt{testa}\sqsubseteq\texttt{plant\ embryo}$). We classify ontological reasoning as the ability to infer such implicit knowledge through reasoning process such as logical deduction.
@@
This capability represents the highest level in our taxonomy and is placed at the top of the triangle in Figure 2. It primarily concerns the process of constructing ontologies. Traditional ontology learning tasks have largely focused on generating hierarchical structures, while often neglecting other essential structural components like properties and instances. Therefore, we propose that ontological learning should encompass multiple dimensions: the generation of class definitions, the construction of class hierarchies, and the integration of properties and their constraints. In addition, ontology alignment—ensuring consistency across multiple ontologies by identifying and mapping equivalent concepts—is a critical aspect and is thus also considered part of this capability.
@@ 3.2 Data Collection and Processing
OntoURL draws on 40 expert-created, open-source ontologies spanning a broad range of 8 different domains, including (1) sciences; (2) health and medicine; (3) business and finance; (4) earth and environment; (5) arts, media and entertainment; (6) food and agriculture; (7) human and society; and (8) the legal domain. All ontologies are provided in RDF Miller (1998) or OWL McGuinness et al. (2004) format. While open-domain ontologies such as DBpedia Auer et al. (2007) are available, we focus on domain-specific ontologies due to their greater depth, consistency, and formal structure. In cases where multiple ontologies exist within the same domain, we address their heterogeneity by designing prompts tailored to each ontology.
@@
After collecting the ontologies, we apply a four-step pipeline to create multiple-choice, true/false question and open-ended questions, as illustrated in Figure 3.
First, we extract task-relevant entities, including classes, instances, properties and their associated semantic details (e.g., definitions, relationships, and range) from each ontology. The sub-ontology of “Basenji” in the first step of Figure 3 presents the most involved elements in this extraction process. Particularly, an ontology reasoner is required to derive implicit relations for reasoning (e.g. the relation between “Basenji” and “Canid”).
@@
After question generation, we apply several filtering and balancing strategies to ensure quality and diversity: (a) To avoid over-representing abstract concepts, we assign sampling probabilities based on class depth (distance from the root) and sample data according to these probabilities. (b) Questions for the same classes are de-duplicated. For instance, as shown in Figure 3, a class like “Dog” has multiple subclasses (e.g., “Basenji”, “Leonberger”), but only one question about the subsumptions of "Dog" will be retained. (c) We limit the number of questions per class (they may from different ontologies) to a maximum of five per task. (d) We set the total number of questions per ontology-task pair at 500 to prevent any single ontology from disproportionately influencing the evaluation results.
Finally, we perform an human verification to ensure the data quality in three key dimensions: (1) Syntax, to verify fluentness and grammatical correctness; (2) Semantics, to confirm the correctness of the answer; and (3) Quality (for multiple-choice questions), to assess whether the distractors are well-designed and appropriately challenging. Each expert reviewed 20% of the data for each task, with a 5% overlap between annotators to allow cross-verification of annotation consistency. This confirms the high quality of the benchmark, with over 95% of questions rated as syntactically and semantically correct, and inter-annotator agreement exceeding 0.85 (Fleiss’ $\kappa$, Landis and Koch, 1977) across evaluation dimensions (detailed guidelines and results are provided in Appendix C).